Required information

[The following information applies to the questions displayed below.]

Fiber Technology, Inc., manufactures glass fibers used in the communications industry. The company's materials and parts manager is currently revising the inventory policy for XL-20, one of the chemicals used in the production process. The chemical is purchased in 10-pound canisters for $96 each. The firm uses 4,000 canisters per year. The controller estimates that it costs $151 to place and receive a typical order of XL-20. The annual cost of storing XL-20 is $4.10 per canister.

Required:

  1. Use the EOQ formula to determine the optimal order quantity.

  2. What is the total annual cost of ordering and storing XL-20 at the economic order quantity?

  3. How many orders will be placed per year?

  4. Fiber Technology’s controller, Jay Turnbull, recently attended a seminar on JIT purchasing. Afterward he analyzed the cost of storing XL-20, including the costs of wasted space and inefficiency. He was shocked when he concluded that the real annual holding cost was $20.20 per canister. Turnbull then met with Doug Kaplan, Fiber Technology’s purchasing manager. Together they contacted Reno Industries, the supplier of XL-20, about a JIT purchasing arrangement. After some discussion and negotiation, Kaplan concluded that the cost of placing an order for XL-20 could be reduced to just $21. Using these new cost estimates, Turnbull computed the new EOQ for XL-20.

  1. a. Use the equation approach to compute the new EOQ.

  2. b. How many orders will be placed per year?
Answer #1

EOQ :-

   √(2)[(annual requirement)(Cost per order)] / annual carrying cost per unit

√(2)(4000)($151) / $4.10

√294634.15

= 542.80 or 543

Total Annual cost of ordering & storing :-

[(annual requirement/EOQ) * Cost per order] + [(EOQ/2) * annual carrying cost per unit]

[(4000/543) * $151] + [(543/2) * $4.10

=$2226

Orders placed =   Orders = Annual requirements/EOQ

    = 4000/543 = 7.37 or 8 orders

After new Estimation (JIT)

(a) EOQ:-

   √(2)[(annual requirement)(Cost per order)] / annual carrying cost per unit

√(2)(4000)($21) / $20.20

91.20 or 92

(b) How many orders?

   Orders = Annual requirements/EOQ

    = 4000/92 = 43.47 or 44 orders